The energy given off as heat by 300g of an alloy as it cools through 50OC raises the

temperature of 300g of water from 30OC to 40 OC. What is the specific heat of the alloy (in

cal/g· OC)?


1
Expert's answer
2021-02-04T19:09:25-0500
"m_{alloy}c_{alloy}t_{alloy}=m_{water}c_{water}(t_{water,f}-t_{water,i}),""c_{alloy}=\dfrac{m_{water}c_{water}(t_{water,f}-t_{water,i})}{m_{alloy}t_{alloy}},""c_{alloy}=\dfrac{300\ g\cdot1.0\ \dfrac{cal}{g\cdot\!^{\circ}C}\cdot(40^{\circ}C-30^{\circ}C)}{300\ g\cdot50^{\circ}C}=0.2\ \dfrac{cal}{g\cdot\!^{\circ}C}."